Find (f of g) (3) when f(x)= 3x^2-2x+1 and g(x)=sqrt(5x-6)
sqrt: means square root

Answer by stanbon(75887)

Find (f of g) (3) when f(x)= 3x^2-2x+1 and g(x)=sqrt(5x-6)
---
fog(3) = f[g(3)]
= f[sqrt(15-6)]
= f[sqrt(9)]
= f[3]
= 3(3^2)-2(3)+1
= 27-6+1
= 22
==============
Cheers,
Stan H.
==============